OpenStudy (anonymous):

Solve 3c + 2 = 4c - 9.

OpenStudy (anonymous):

3c + 2 = 4c - 9. 3c+2 - 3c = 4c - 9 - 3c 2 = c - 9 2 + 9 = c - 9 +9 c = 11

OpenStudy (anonymous):

4x - 2x - 2 = -42. Here is another one.

OpenStudy (anonymous):

try it just as above i did

OpenStudy (anonymous):

Is it a negative or positive.

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

4x - 2x - 2 = -42. 2x - 2 = -42 2x - 2 +2 = -42 +2 2x = -40 2x/2 = -40/2 x=-20

OpenStudy (anonymous):

Okay, I see it but i don't get how you got that. I got negative 7.

OpenStudy (anonymous):

Never mind I see you added 4x to - 2x

OpenStudy (anonymous):

okay finally u got it right???

OpenStudy (anonymous):

Yes. Help with this one.\[|3u - 18| = 0\]

OpenStudy (anonymous):

(6,-6)

OpenStudy (anonymous):

No just 6.

OpenStudy (anonymous):

Two consecutive odd integers have a sum of 48. What is the largest of the two integers?

OpenStudy (anonymous):

23+25 = 48 so 25